[Brahms-l] Shift report 20050603 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Fri Jun 03 2005 - 23:52:18 EDT
Shifter leader was HHD
Shifters were 


1. MCR took their time establishing the beam; but once it was up there were no problems.

2. Data runs:
   FS@4B2442, MRS@60A700: 15167*,15168*,15169*,15170
   TPC pedestal run: 15165
   non-TPC pedestal run: 15166

3. The BB voltages are jamming. There was a problem with the TPC voltages being off (the runs marked with * above were taken during this) but JHL fixed it.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Jun 3 23:52:33 2005

This archive was generated by hypermail 2.1.8 : Fri Jun 03 2005 - 23:52:46 EDT